A Quick Guide to the Technician Class HAM Exam

Derek A. Atkins

N1NWH January 16, 1993

What is a HAM?

Well, if you don’t know what a HAM is, you probably are reading the wrong document. A HAM is a nickname for an Amateur Radio operator. This document will explain to you the basics of what is on the exam, things you should know, and how to study and pass the Technician Class amateur radio exam.

What, Where, When, and How?

The next MIT Ham exam is at 7:30pm on Wednesday, January 20, 1993. The Technician Class license test is comprised of two exams, a 30 question novice exam (part 2) and a 25 question technician exam (part 3A). Both exams are multiple choice. You must have at least 75% of the questions correct to pass, and the exams must be taken, and passed, sequentially (this means you must pass the novice exam before you can take the technician exam). The cost of the exam is currently $5.40, and you can take as many exams as you want, until you fail! There is also the Morse Code exam. To get your Technician Plus, you need to pass at least the 5 wpm Morse Code Exam (part 1A).

The Important Stuff

The following sections contain most of the stuff, which you probably don’t know, that you should know for the exam. I won’t go into the technical details, since I am assuming you’ve taken the equivalent of 6.002 and 6.003. If you haven’t, please feel free to ask questions. The FCC question pools are online on Athena. You can get to them from any Athena workstation by executing the command:

   cd /afs/sipb.mit.edu/project/ham/questions
   more <filename>

where <filename> is the question pool you want to read. In that directory are the question pools for all the ham classes. There are is also other information available under the ham directory, including the FCC Rules (part 97).

Frequencies

You should memorize tables of frequency bands available. There are other bands allocated to Amateur radio than those listed in these tables, however I didn’t list those without any privileges to conserve space. Even though not all of it applies to you, as a Technician, you need to know it because it will be on the exams, assuredly. These are listed in tables 1 and 2.

There are a couple of data points between Novice and Tech Plus that should be pointed out since there are questions on them. For the Novice Class License, the operating privileges in the 220 MHz band is limited to 222.1–223.91 MHz, and in the 1270 MHz band the limits are 1270–1295 MHz. Remember that the Novice has no operating privileges in the 6 Meter, 2 Meter, and 440 MHz bands! Another point of information is that Technicians can send phone transmissions in the 6 Meter band on 50.1–54.0 MHz, and FM image transmissions in the 2 Meter band on 144.1–148.0 MHz. Be sure to read over the question pools!

Other questions ask about wavelengths for a frequency. If you remember this shortcut you can’t go wrong.

λ(m) = \frac{300}{f(MHz)}

Also remember that the length of a half-wave dipole antenna can be calculated using the formula

length(m) = \frac{468}{f(MHz)}

Call Signs

Every HAM is issued a Call Sign which uniquely identifies them in the world! A Call Sign must have one or two letters preceding a number, and then one, two, or three letters after the number. The first characters of the call sign identify the country and/or geographic region where the call sign was originally issued. In the US, all call signs start with the letters A, K, N, or W. Each station must identify itself at least once every ten minutes during a conversation and at the end of the transmission.

Antennas

HAMs need to know about RF propagation and antennas. We won’t go into a lot of it here, since what you really need to know are the answers to the questions. I will say that you should know a little about the Yagi antenna, which are a bunch of half-wavelength elements with parasitically driven parallel side elements. Another is the cubical-quad antenna, which is a square full-wave closed loop (each side is 1/4 λ) with parallel parasitic elements. There are a number of questions about antennas on the exams, and you should read them over and learn them.

You should also know about Standing Wave Ratios, or SWRs. Remember that you want an SWR close to unity! A high SWR is bad, and means that something is wrong; either the antenna is the wrong length or there is a short. Another thing to remember is that if the SWR increases as you increase frequency, the antenna is too long, and if it decreases with an increase in frequency it’s too short.

Other random things to know

The questions on the exams basically are a lot of memorization or technical questions. I expect that most MIT students who have had any course VI experience can answer most of the technical questions.

An example of things you should know is the five principles: Recognition of emergency communications, advancement of the radio art, improvement of communication and technical skills, increase in the number of trained radio operators and electronics experts, and the enhancement of international goodwill. MEMORIZE these! It WILL be on the exam! These are the fundamental purpose of the amateur service.

Another piece of random data that you should probably know is the effect of sunspots on radio communication. Remember that an average sunspot cycle is 11 years long, and that the more sunspots there are, the higher the maximum usable frequency. This also means that more sunspots means a higher level of ionization in the atmosphere.

You should know about interference: different causes, possible effects, and what to do if you get interference or someone is getting interference from you. All of these are in the question pools. Read them over.

You should also know what the different types of modulation identifiers mean. For example, A1A is CW modulation, or F3B is FM. The first letter is modulation type, the second is the nature of the modulating signal, and the third is the type of information.

If I haven’t said it enough, I’ll say it again! Read the question pools. Read through at least once, try to answer all the questions. The answers are at the end to the files. These are the exact questions that will appear on your exam. If you learn the questions, you will have an easier time on the exam, since you will have at least seen them once before. Remember, the exam is on Wednesday, so you have 4 days to prepare! This is plenty of time to learn the material and pass the exams!

73’s N1NWH

Band  Frequency  Emissions Comments
6 Meters  50.0–54.0 MHz All modes Sideband voice, radio control, FM/FM

repeater, digital computer, remote bases, and autopatches, Even CW. (1500 watts PEP output limitation)

2 Meters 144–148 MHz All modes All types of operation including satellite

and owning repeater and remote bases. (1500 watts PEP output limitation)

11/4 Meters 222–225 MHz  All modes (1500 watt PEP output limitation)
70 cm 420–450 MHz All modes Including amateur television, packet,

RTTY, FAX, and FM voice repeaters (1500 watt PEP limit)

35 cm 902–928 MHz All modes Plenty of room! (1500 watts PEP limit)
23 cm  1240–1300 MHz All modes All privileges

Table 1: Technician Class Operating Privileges

Band  Frequency  Emissions  Comments
80 Meters  3675–3725 KHz  Code Only  **Limited to Morse code

(200 watt PEP output limitation)

40 Meters  7100–7150 KHz  Code Only  (200 watt PEP limit)
15 Meters  21100–21200 KHz Code Only  (200 watt PEP limit)
10 Meters  28100–28500 KHz Code  Morse code and digital computer
28100–28300 KHz Data and code (200 watt PEP output code limitation)
28300–28500 KHz Phone and code Sideband voice (200 watt PEP limit)
6 Meters 50.0–54.0 MHz All modes Sideband voice, radio control, FM/FM

repeater, digital computer, remote bases, and autopatches, Even CW. (1500 watts PEP output limitation)

2 Meters 144–148 MHz All modes All types of operation including satellite

and owning repeater and remote bases. (1500 watts PEP output limitation)

11/4 Meters 222–225 MHz  All modes  (1500 watt PEP output limitation)
70 cm  420–450 MHz  All modes  Including amateur television, packet,

RTTY, FAX, and FM voice repeaters (1500 watt PEP limit)

35 cm  902–928 MHz  All modes  Plenty of room! (1500 watts PEP limit)
23 cm  1240–1300 MHz  All modes  All privileges

Table 2: Technician Plus Class Operating Privileges

A Alfa  F Foxtrot K Kilo  P Papa  U Uniform Z Zulu
B Bravo G Golf  L Lima  Q Quebec V Victor
C Charlie H Hotel  M Mike  R Romeo W Whiskey
D Delta I India  N November S Sierra X X-ray
E Echo  J Juliet O Oscar  T Tango Y Yankee

Table 3: Phonetic Alphabet

Readability  Signal Strength  Tone
1 Unreadable  Faint and barely  Very rough, broad signals,
perceptible signals  60 cycle AC may be present
2 Barely readable; occasional Very weak signals  Very rough AC tone,
words distinguishable.  harsh, broad
3 Readable with considerable Weak signals  Rough, low pitched AC tone,
difficulty.  no filtering
4 Readable with practically Fair signals  Rather rough AC tone,
no difficulty  some trace of filtering
5 Perfectly readable Fairly good signals Filtered rectified AC note,

musical, ripple modulated

6 Good signals Slight trace of filtered tone but

with ripple modulation

7  Moderately strong signals Near DC tone but trace of

ripple modulation

8 Strong signals Good DC tone, may have slight trace

of modulation

9  Extremely strong signals Purest, perfect DC tone with no trace

of ripple modulation

Table 4: RST Reporting Chart (Tone is only for CW)

QRL Are you busy? (I am busy with ) Please don’t interfere.
QRS Shall I send more slowly? Send more slowly ( WPM)
QRT Shall I stop sending? Stop sending
QRU Have you anything for me? I have nothing for you.
QRV Are you ready? I am ready.
QRZ Who is calling me? You are being called by (on KHz)
QSL Can you acknowledge receipt? I am acknowledging receipt.
QSO Can you communicate with  direct or by relay? I can ...
QTH What is your location? My location is
QTR What is the time? The time is

Table 5: A partial list of Q-codes